OpenStudy (anonymous):

Hi! I would like to ask help on solving this: 400000=600000(1+r)^5. Your help would be greatly appreciated. Thanks ahead!

OpenStudy (anonymous):

First divide by 600000, then take the 5th root of both sides to eliminate the ^5. Since its an odd exponent its ok to have a negative under the radical. Then you'll have some number (-20000^(1/5))=(1+r). Subtract one from both sides to find r.

OpenStudy (anonymous):

I'm sorry. If I'm not mistaken, you mean I should divide both sides by 600000 right? So 400000/600000=0.666. May I ask how you were able to arrive to -20000?

OpenStudy (anonymous):

just noticed that that 6 number had one more 0 than the 4 one haha so my number is wrong. But yea divide. I didn't notice that 0 so you're right, you'd get 2/3.

OpenStudy (anonymous):

just ignore me...my math is very off. I subtracted. But yes, you divide and you'd get what you got.

OpenStudy (anonymous):

Yeah. That's what I actually did at first. Just wanted to confirm it 'cause when I checked it, it seems to be wrong. O.o

OpenStudy (anonymous):

hmm..well the 5th root of that will be pretty ugly...but its the only way I can see to possibly do that

OpenStudy (anonymous):

Definitely ugly. Let me check again. Thanks for the help, though! :D

OpenStudy (anonymous):

Gosh, it's waaaaaaaaaaaaaay too far. :/

OpenStudy (anonymous):

\[\sqrt[5]{2}\] \[\sqrt[5]{3}\] If those were over each other as a fraction, you'd have to rationalize ( I think thats the word?) so you don't have a sqrt in the denominator. Uh oh...way too far?

OpenStudy (anonymous):

Yeah. Haha. The answer didn't even reach like beyond 50.

OpenStudy (anonymous):

Oh this shouldn't even be 1 I don't think?

OpenStudy (anonymous):

Actually, I'm not also so sure with the equation. This is the problem: A new car that costs Php 600 000 decreased in value to Php 400 000 after 5 years. Find the average annual rate of depreciation.

OpenStudy (anonymous):

huh..well I have no idea what you'd do for that. My only understanding of depreciation comes from basic accounting..

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

Hmm. 25 something hahaah

OpenStudy (anonymous):

I'll try to use another formula.

OpenStudy (anonymous):

yea that isn't the right formula, it looks like its the formula for annual interest if I'm not mistaken

OpenStudy (anonymous):

I think it is. I just tried it out if it would work. haha

OpenStudy (anonymous):

haha yea not so much!

OpenStudy (anonymous):

Hey, it's strange. I used a formula for depreciation rate and I arrived at the same result.

OpenStudy (anonymous):

haha that is strange

OpenStudy (anonymous):

:(((((